# Graph Masked Autoencoder for Spatio-Temporal Graph Learning

### v1

@@ IV-G Model Efficiency Study (RQ5)
We investigate the efficiency of our method STGMAE and existing region representation learning methods  w.r.t training time. We provide results in Table V. All of these methods are adopted in python 3.8, pytorch 1.7.0 with cuda 11.3 and tensorflow 1.15.3 (GPU version). The experiments are conducted in the same server. From the table, By comparing the model efficiency of STGMAE to the baselines (e.g., HDGE, MVURE and MGFN), we can see that it is competitive, indicating that it has the capacity to handle massive amounts of spatial-temporal data. The performance of the region representation is greatly improved by the generative self-supervised learning paradigms while incurring little additional processing expense.
@@
Table V: Computational cost (seconds) of our STGMAE and SOTA spatial-temporal region representation methods. MAE and MAPE are estimated for NYC crime prediction results.
@@
Figure 6: Case study of STGMAE on New York City. Our STGMAE can alleviate the connection bias in spatial-temporal graphs by refining the salient relational signals between far away regions (region 1 & region 124).
IV-H Case Study (RQ6)
In this part, we use a case study to demonstrate how well our STGMAE can understand the relationship between global regions and geographical semantics in Figure 6. We sample 1 region pairings, or distant regions, specifically (region-1, region-124). Even though regions 1 and 158 are geographically far apart, as seen in the provided Figure, their latent meanings are well preserved in our encoded embeddings. Their cluster results from the MGFN baseline, however, demonstrate a significant difference. In contrast, the region clusters discovered using our STGMAE method exhibit similarity. These findings show that our STGMAE has the advantage of incorporating global region dependencies over the entire metropolitan landscape, which is important in real-life applications.
V Conclusion
@@
We introduce a novel generative spatial-temporal graph self-supervised learning paradigm for region representation. Our approach incorporates mask mechanisms to address limitations in existing area representation methods. Through extensive experiments on real-world datasets and three spatial-temporal mining tasks, we demonstrate the effectiveness of our STGMAE in learning area graph representations from multi-view spatial-temporal data. Our contributions include proposing a self-supervised learning paradigm using generative models and masks, addressing challenges in area representation methods, and achieving superior performance compared to state-of-the-art approaches in various spatial-temporal mining tasks, such as crime prediction.
VI Acknowledgement
@@
This project is partially supported by HKU-SCF FinTech Academy and Shenzhen-Hong Kong-Macao Science and Technology Plan Project (Category C Project: SGDX202108 23103537030) and Theme-based Research Scheme T35-710/20-R. We appreciate the help from Dr. Huang and Dr. Xia.

### v2

@@ 3.7. Model Efficiency Study (RQ5)
We investigate the efficiency of our method HGAurban and existing region representation learning methods  w.r.t training time. We provide results in Table 4. All of these methods are adopted in python 3.8, pytorch 1.7.0 with cuda 11.3 and tensorflow 1.15.3 (GPU version). The experiments are conducted in the same server. From the table, By comparing the model efficiency of HGAurban to the baselines (e.g., HDGE, MVURE and MGFN), we can see that it is competitive, indicating that it has the capacity to handle massive amounts of spatial-temporal data. The performance of the region representation is greatly improved by the generative self-supervised learning paradigms while incurring little additional expense.
@@
Table 4. Computational cost (seconds) of our HGAurban and SOTA spatial-temporal region representation methods. MAE and MAPE are estimated for NYC crime prediction results.
@@
4. Related Work
Self-supervised Learning for Spatial-Temporal Graphs. In the domain of self-supervised learning (SSL) for spatial-temporal graphs, a line of research has emerged, focusing on graph-based SSL methods specifically designed for capturing the spatial and temporal dependencies within the graph structure. Several recent studies have contributed to this area, including  (Liu et al., 2023; Ji et al., 2023; Zhang et al., 2023a, b), and  (Ji et al., 2022). Liu et al. (Liu et al., 2023) propose an innovative approach that adopts the self-supervised learning paradigm to leverage supplementary information that is typically inaccessible due to privacy concerns. Ji et al. (Ji et al., 2023) introduce self-supervised learning paradigms in a spatial-temporal traffic prediction framework. The proposed method enhances the traffic pattern representations to capture both spatial and temporal heterogeneity within the spatial-temporal graph, leading to improved traffic prediction performance. Overall, these studies demonstrate the effectiveness of methods of self-supervised learning in capturing spatial and temporal dependencies within spatial-temporal graphs. By leveraging additional information or incorporating self-supervised learning paradigms, these approaches enhance representation learning and improve the performance of downstream tasks in various domains, including POI recommendation, traffic prediction, and region representations on the spatial-temporal graph for downstream tasks.
Graph Mask Autoencoder. Graph autoencoder methods have emerged as a category of generative self-supervised learning approaches that aim to construct input graphs directly, eliminating the need for high-quality data augmentation typically required in contrastive self-supervised learning. An earlier study by Kipf and Welling (Kipf and Welling, 2016) introduces Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE), which utilize Graph Convolutional Networks (GCN) as the encoder and a dot-product decoder. These models focus on reconstructing the topological structure and node content of the graph from a compressed latent representation. ARVGA (Adversarially Regularized Variational Graph Autoencoder) (Pan et al., 2018) proposes a novel adversarial graph embedding framework. It trains a decoder to reconstruct both the topological structure and node content of the graph from a compact representation generated by the framework. This adversarial training enhances the robustness and generalization capabilities of the model. MGAE (Multi-modal Graph Autoencoder)(Wang et al., 2017) and GALA (Graph Autoencoder with a Ladder-like Architecture)(Park et al., 2019) are methods that combine structure and node reconstruction. They aim to learn a comprehensive representation of the graph by simultaneously reconstructing the graph structure and node attributes. In a recent study by Hou et al. (Hou et al., 2022), a graph autoencoder approach called GraphMAE is proposed, which focuses solely on node reconstruction for general graphs. By reconstructing nodes, GraphMAE achieves promising results in node classification tasks.
Region Representation Learning. Research on the region representation learning problem includes the following studies (Wang and Li, 2017; Yao et al., 2018; Zhang et al., 2021, 2019; Fu et al., 2019; Wu et al., 2022; Zhang et al., 2023a). In particular, Fu et al. (Fu et al., 2019) suggest to incorporate both intra-region information (such as POI distance inside a region) and inter-region information (such as the similarity of POI distributions between two regions) to increase the quality of region representations. The concept of putting forward in (Fu et al., 2019) is expanded upon by Zhang et al. (Zhang et al., 2019) by using a collective adversarial training.
In a recent study conducted by Zhang et al. (Zhang et al., 2021), a multi-view joint learning model is developed to learn region representations. To capture region correlations from different perspectives, such as human mobility and region properties, the study incorporates a graph attention technique for each view. However, this approach focuses solely on modeling region correlations and does not consider the inclusion of point-of-interest (POI) data, which is essential for capturing region functionality. Similarly, Wu et al. (Wu et al., 2022) propose a method for learning area representations by extracting traffic patterns. However, their approach also excludes POI data and relies solely on mobility data. The efficacy of these techniques heavily relies on the generation of high-quality region graphs. Furthermore, they may struggle to learn accurate region representations when faced with noisy and skewed spatio-temporal data. In contrast, a recent study by Zhang et al. (Zhang et al., 2023a) proposes the adoption of a contrastive self-supervised learning paradigm on a multi-view heterogeneous spatio-temporal graph. This approach utilizes a self-contrastive learning method to obtain high-quality region vectors. By contrast, our paper employs a generative self-supervised learning method to model region correlations by masking links and nodes on the heterogeneous spatio-temporal graph.
5. Conclusion
@@
We introduce a novel generative spatial-temporal graph self-supervised learning paradigm for region representation. Our approach incorporates mask mechanisms to address limitations in existing area representation methods. Through extensive experiments on real-world datasets and three spatial-temporal mining tasks, we demonstrate the effectiveness of our HGAurban in learning area graph representations from multi-view spatial-temporal data. Our contributions include proposing a self-supervised learning paradigm using generative models and masks, addressing challenges in area representation methods, and achieving superior performance compared to state-of-the-art approaches in various spatial-temporal mining tasks, such as crime prediction.
Acknowledgement
@@
The Australian Research Council partially supports this work under the streams of Future Fellowship (Grant No. FT210100624), the Discovery Project (Grant No. DP240101108), and the Linkage Projects (Grant No. LP230200892 and LP240200546). This project is also partially supported by Shenzhen-Hong Kong-Macao Science and Technology Plan Project (Category C Project: SGDX202108 23103537030) and Theme-based Research Scheme T35-710/20-R. We appreciate help of Dr. Huang and Dr. Xia.